# Effective Polaron Dynamics of an Impurity Particle Interacting with a Fermi Gas

### v1

@@ 3.2 Effective coherent state
$\widetilde{\text{Error}}$ $\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\left(\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)-e^{is\epsilon_{\alpha}(k)}\right)\overline{(h_{y})_{\alpha}(k)}W(\eta_{s})\mathcal{R}_{\alpha}(k)\psi$
$\hphantom{\coloneqq}-\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}W(\eta_{s})\mathcal{R}_{\alpha}(k)^{\ast}(h_{y})_{\alpha}(k)\psi$
$\hphantom{\coloneqq}-2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{k,\alpha}W_{\tau}(\eta_{s})\psi$ (6.5)
@@
$\sup_{s\in[0,t]}\left\{\|\widetilde{\text{Error}}\|+\|h_{0}W(\eta_{s})\psi\|\right\}$
$\leq C\lambda k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta}(f_{C}(\kappa k_{\textnormal{F}}t)-1)+C\beta\left(\log\left[f_{1}(\kappa k_{\textnormal{F}}t)\right]+\log\left[f_{1}(\kappa k_{\textnormal{F}}t)\right]^{4}\right)\left(f_{C}(\kappa k_{\textnormal{F}}t)+1\right)$
$\leq CQ_{V}(\kappa k_{\textnormal{F}}t)\max\{\lambda k_{\textnormal{F}}MN^{-\frac{2}{3}+\delta},\beta\}\eqqcolon d\in\mathcal{O}(\max\{\lambda k_{\textnormal{F}}^{\epsilon},\beta\})$ (6.10)
using that for all $t\in[0,T]$ with $T\in\mathcal{O}(\tau)$ it holds $\kappa k_{\textnormal{F}}t\in\mathcal{O}(1)$.
@@
$\leq\|h_{y}\|\|(\mathcal{N}+1)^{1/2}\psi\|+2\sup_{s\in[0,t]}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\frac{1-\cos(s\epsilon_{\alpha}(k))}{\epsilon_{\alpha}(k)}|(h_{y})_{\alpha}(k)|^{2}$ (6.11)
$\leq\|h_{y}\|+\frac{2}{\kappa k_{\textnormal{F}}}\|h_{y}\|^{2}\leq C\lambda k_{\textnormal{F}}\big{(}\|\hat{V}\|+\kappa^{-1}\lambda\|\hat{V}\|^{2}\big{)}\eqqcolon\alpha.$ (6.12)
@@
$\left|\int_{0}^{t}\textnormal{d}s\,\text{Re}\langle h_{y},\eta_{s}\rangle\right|$ $=\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}|(h_{y})_{\alpha}(k)|^{2}\right|$
$=\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma^{\textnormal{nor}}}\lambda^{2}\hat{V}(k)^{2}\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}n_{\alpha}(k)^{2}\right|$
$=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{\kappa k_{\textnormal{F}}}\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}\frac{e^{-i2\kappa k_{\textnormal{F}}|k|s}-1}{2\kappa k_{\textnormal{F}}|k|s}\right|\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$ (6.13)
@@
$\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}n_{\alpha}(k)^{2}$ $=2\pi\int_{0}^{\pi/2}\textnormal{d}\theta\ \frac{e^{-is2\kappa k_{\textnormal{F}}|k|\cos\theta}-1}{2\kappa k_{\textnormal{F}}|k|\cos\theta}\cos\theta\sin\theta\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$
$=i\frac{2\pi k_{\textnormal{F}}^{2}}{2\kappa k_{\textnormal{F}}}\frac{e^{-i2\kappa k_{\textnormal{F}}|k|s}-1}{2\kappa k_{\textnormal{F}}|k|s}\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}.$ (6.14)
@@
$\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}\frac{e^{-i2\kappa k_{\textnormal{F}}|k|s}-1}{2\kappa k_{\textnormal{F}}|k|s}\right|=\left|\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}\int_{0}^{t}\textnormal{d}s\frac{1-\cos(2\kappa k_{\textnormal{F}}|k|s)}{2\kappa k_{\textnormal{F}}|k|s}\right|.$ (6.15)
Note that it is well-known that for all $x>0$ it holds $\cos(x)<1-4x^{2}/\pi^{2}$ and thus $1-\cos(x)\geq 4x^{2}/\pi^{2}$ for all $x\in(0,\pi/2)$ Also it holds for all $x>\pi/2$ that
@@
$\geq\frac{2\pi\lambda^{2}k_{\textnormal{F}}^{2}}{\kappa k_{\textnormal{F}}}\sum_{k\in\Gamma^{\textnormal{nor}}}\frac{\hat{V}(k)^{2}}{2\kappa k_{\textnormal{F}}|k|}\bigg{(}\chi(2\kappa k_{\textnormal{F}}|k|t>\frac{\pi}{2})\left\{\frac{1}{2}+\ln(2\kappa k_{\textnormal{F}}|k|t)-\ln(\frac{\pi}{2})\right\}$
$+\chi(2\kappa k_{\textnormal{F}}|k|t\leq\frac{\pi}{2})\frac{2(2\kappa k_{\textnormal{F}}|k|t)^{2}}{\pi^{2}}\bigg{)}\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$
$\eqqcolon b_{t}+td$ (6.18)
with $d\in\mathcal{O}(MN^{-\frac{2}{3}+\delta})$ defined in (6.10) and
$\dot{b}_{t}$ $=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{\kappa k_{\textnormal{F}}}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}\left(\chi(2\kappa k_{\textnormal{F}}|k|t>\frac{\pi}{2})\frac{1}{2\kappa k_{\textnormal{F}}|k|t}+\chi(2\kappa k_{\textnormal{F}}|k|t\leq\frac{\pi}{2})\frac{8\kappa k_{\textnormal{F}}|k|t}{\pi^{2}}\right)-d>-d.$ (6.19)
@@
$g_{t}\geq b_{t}-\alpha\int_{0}^{t}\textnormal{d}sg_{s}$ (6.20)
with $b_{t}$ and $\alpha$ defined in (6.18) and (6.12), respectively.
@@
$h_{t}$ $=b_{t}-\alpha\int_{0}^{t}\textnormal{d}sh_{s}\qquad\text{with }h_{0}=0.$ (6.21)
@@
Assume for the proof by contradiction $\exists t_{0}>0:h_{t_{0}}>g_{t_{0}}$ and set $I\subseteq\mathbb{R}_{>0}$ as largest open interval satisfying $t_{0}\in I$ and $h_{t}>g_{t}$ for all $t\in I$. Note that since $h_{t}$ is differentiable such an open interval exists. It holds $t_{1}\coloneqq\inf I>0$ since $h_{0}=0=g_{0}$ and for all $t\in I$
$\dot{g}_{t}\geq\dot{b}_{t}-\alpha g_{t}>\dot{b}_{t}-\alpha h_{t}=\dot{h}_{t}$
@@
Thus we obtain the desired contradiction to $t_{1}$ being the inf of the largest set satisfying $h_{t}>g_{t}$. ∎
@@
$h_{t}=e^{-\alpha t}\int_{0}^{t}\dot{b}_{s}e^{\alpha s}\textnormal{d}s.$ (6.22)
@@
$h_{t}$ $=e^{-\alpha t}\int_{0}^{t}\dot{b}_{s}e^{\alpha s}\textnormal{d}s$
$=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{\kappa k_{\textnormal{F}}}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}e^{-\alpha t}\int_{0}^{t}\bigg{(}\chi(2\kappa k_{\textnormal{F}}|k|s>\frac{\pi}{2})e^{\alpha s}\frac{1}{2\kappa k_{\textnormal{F}}|k|s}$
$+\chi(2\kappa k_{\textnormal{F}}|k|s\leq\frac{\pi}{2})e^{\alpha s}\frac{8\kappa k_{\textnormal{F}}|k|s}{\pi^{2}}\bigg{)}\textnormal{d}s+d(e^{-\alpha t}-1)$ (6.23)
$\geq\frac{\pi\lambda^{2}k_{\textnormal{F}}}{\kappa}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}\bigg{\{}\chi(|k|t\leq\frac{\pi}{4\kappa k_{\textnormal{F}}})\frac{8\kappa k_{\textnormal{F}}|k|}{\alpha^{2}\pi^{2}}\left(e^{-\alpha t}+\alpha t-1\right)+$
$+\chi(|k|t>\frac{\pi}{4\kappa k_{\textnormal{F}}})\frac{8\kappa k_{\textnormal{F}}|k|}{\alpha^{2}\pi^{2}}\big{(}e^{-\alpha\pi/(4\kappa k_{\textnormal{F}}|k|)}+\frac{\alpha\pi}{4\kappa k_{\textnormal{F}}|k|}-1\big{)}\bigg{\}}+d(e^{-\alpha t}-1)$ (6.24)
$\geq\frac{\lambda^{2}k_{\textnormal{F}}^{2}}{\alpha^{2}\pi}\sum_{k\in\Gamma^{\textnormal{nor}}}\hat{V}(k)^{2}|k|\min\bigg{\{}f(t),f\big{(}\frac{\pi}{4\kappa k_{\textnormal{F}}|k|}\big{)}\bigg{\}}+d(e^{-\alpha t}-1)$ (6.25)
with $f(t)\coloneqq\left(e^{-\alpha t}+\alpha t-1\right)$ defining a non-negative monotonically increasing function. Since $d\in\mathcal{O}(MN^{-\frac{2}{3}+\delta})$ we obtain the desired result that (5.20) has a lower bound of order 1. ∎
@@ Appendix A Estimates within the bosonization framework
In this section we collect all relevant estimates of the bosonization framework which was first developed in [BNP+19, BNP+21a, BNP+21b, BPSS23] in the semiclassical regime. In addition to the references we provide brief proof sketches in order to stress the dependence of the coupling constants.
@@
For $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta}$ and $k\in\Gamma^{\textnormal{nor}},\alpha\in\mathcal{I}_{k}$ it holds
@@
For $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta}$ and $k\in\Gamma^{\textnormal{nor}}$ it holds
@@
Lemma A.3 (Estimates on the CAR error term, [BNP+21b, Lemma 5.2]).
For $k^{\prime},k\in\Gamma^{\textnormal{nor}}$ and $\alpha\in\mathcal{I}_{k},\beta\in\mathcal{I}_{k^{\prime}}$ the error term $\mathcal{E}_{\alpha}(k,k^{\prime})$ as defined in (2.22) satisfies $\mathcal{E}_{\alpha}(k,k^{\prime})=\mathcal{E}_{\alpha}(k^{\prime},k)^{\ast}$, commutes with $\mathcal{N}$ and for all $\gamma\in\mathcal{I}_{k}\cap\mathcal{I}_{k^{\prime}}$ it holds for all $\zeta\in\mathcal{F}$
@@
It holds for all $k\in\Gamma^{\textnormal{nor}}$ and $\psi\in\mathcal{F},f\in l^{2}(\mathcal{I}_{k})$:
@@
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}(k)\psi\|^{2}\leq\|\mathcal{N}^{\frac{1}{2}}\psi\|^{2}$
@@
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}(k)\psi\|\leq M^{\frac{1}{2}}\|\mathcal{N}^{\frac{1}{2}}\psi\|$
@@
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)\psi\|\leq M^{\frac{1}{2}}\|(\mathcal{N}+M)^{\frac{1}{2}}\psi\|$
@@
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)\psi\|^{2}\leq\|(\mathcal{N}+M)^{\frac{1}{2}}\psi\|^{2}$
@@
    $\|\sum_{\alpha\in\mathcal{I}_{k}}f_{\alpha}c_{\alpha}^{\ast}(k)\psi\|\leq\|f\|_{l^{2}}\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|$
@@
    $\sum_{\alpha\in\mathcal{I}_{k}}c_{\alpha}^{\ast}(k)c_{\alpha}(k)\leq\mathcal{N}$
Lemma A.5 (Error of linearized of kinetic energy, [BNP+21b, Lemma 8.2]).
It holds for $k\in\Gamma^{\textnormal{nor}}$, $\alpha\in\mathcal{I}_{k}$ and all $\psi\in\mathcal{F}$
$[\mathbb{H}_{0},c_{\alpha}^{\ast}(k)]=2\kappa k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)^{\ast}$
@@
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\psi\|^{2}$ $\leq C\left(\kappa N^{\frac{1}{3}}M^{-\frac{1}{2}}\right)^{2}\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|^{2}$
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\psi\|$ $\leq C\kappa N^{\frac{1}{3}}\|\mathcal{N}^{\frac{1}{2}}\psi\|.$

### v2

@@ 3.2 Effective coherent state
$\widetilde{\text{Error}}$ $\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\left(\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)-e^{is\epsilon_{\alpha}(k)}\right)\overline{(h_{y})_{\alpha}(k)}W(\eta_{s})\mathcal{R}^{1}_{\alpha}(k)\psi$
$\hphantom{\coloneqq}-\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}W(\eta_{s})\mathcal{R}^{1}_{\alpha}(k)^{\ast}(h_{y})_{\alpha}(k)\psi$
$\hphantom{\coloneqq}-2i\int_{0}^{1}\textnormal{d}\tau\ W_{1-\tau}(\eta_{s})\text{Im}\langle\dot{\eta}_{s},\mathcal{R}^{1-\tau}\rangle_{\Gamma}W_{\tau}(\eta_{s})\psi.$ (6.5)
@@
$t\sup_{s\in[0,t]}\left\{\|\widetilde{\text{Error}}\|+\|h_{0}W(\eta_{s})\psi\|\right\}$
$\leq C\lambda k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta}(f_{C}(\lambda k_{\textnormal{F}}t)-1)+C\beta\left(\log\left[f_{1}(\lambda k_{\textnormal{F}}t)\right]+\log\left[f_{1}(\lambda k_{\textnormal{F}}t)\right]^{4}\right)\left(f_{C}(\lambda k_{\textnormal{F}}t)+1\right)$
$\leq CQ_{V}(k_{\textnormal{F}}t)\max\{\lambda k_{\textnormal{F}}tMN^{-\frac{2}{3}+\delta},\beta t\}$
$\leq C\max\{MN^{-\frac{2}{3}+\delta},\beta k_{\textnormal{F}}^{-1}\lambda^{-1}\}\eqqcolon d$ (6.10)
where we used $0\leq t\lesssim\mathcal{O}(k_{\textnormal{F}}^{-1}\lambda^{-1})$ and defined the error variable $d$.
@@
$\leq\|h_{y}\|\|(\mathcal{N}+1)^{1/2}\psi\|+2\sup_{s\in[0,t]}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\frac{1-\cos(s\epsilon_{\alpha}(k))}{\epsilon_{\alpha}(k)}|(h_{y})_{\alpha}(k)|^{2}$ (6.11)
$\leq\|h_{y}\|+\frac{2}{k_{\textnormal{F}}}\|h_{y}\|^{2}\leq C\lambda k_{\textnormal{F}}\big{(}\|\hat{V}\|+\lambda\|\hat{V}\|^{2}\big{)}\eqqcolon\theta$ (6.12)
where we introduced the variable $\theta$.
@@
$\left|\int_{0}^{t}\textnormal{d}s\,\text{Re}\langle h_{y},\eta_{s}\rangle\right|$ $=\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}|(h_{y})_{\alpha}(k)|^{2}\right|$
$=\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma}\lambda^{2}\hat{V}(k)^{2}\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}n_{\alpha}(k)^{2}\right|$
$=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{k_{\textnormal{F}}}\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma}\hat{V}(k)^{2}\frac{e^{-i2k_{\textnormal{F}}|k|s}-1}{2k_{\textnormal{F}}|k|s}\right|\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$ (6.13)
@@
$\sum_{\alpha\in\mathcal{I}_{k}}\frac{e^{-is\epsilon_{\alpha}(k)}-1}{\epsilon_{\alpha}(k)}n_{\alpha}(k)^{2}$ $=2\pi\int_{0}^{\pi/2}\textnormal{d}\theta\ \frac{e^{-is2k_{\textnormal{F}}|k|\cos\theta}-1}{2k_{\textnormal{F}}|k|\cos\theta}\cos\theta\sin\theta\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$
$=i\frac{2\pi k_{\textnormal{F}}^{2}}{2k_{\textnormal{F}}}\frac{e^{-i2k_{\textnormal{F}}|k|s}-1}{2k_{\textnormal{F}}|k|s}\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}.$ (6.14)
@@
$\left|\int_{0}^{t}\textnormal{d}s\ \text{Re}\sum_{k\in\Gamma}\hat{V}(k)^{2}\frac{e^{-i2k_{\textnormal{F}}|k|s}-1}{2k_{\textnormal{F}}|k|s}\right|=\left|\sum_{k\in\Gamma}\hat{V}(k)^{2}\int_{0}^{t}\textnormal{d}s\frac{1-\cos(2k_{\textnormal{F}}|k|s)}{2k_{\textnormal{F}}|k|s}\right|.$ (6.15)
It is well-known that for all $x>0$ it holds $\cos(x)<1-4x^{2}/\pi^{2}$ and thus $1-\cos(x)\geq 4x^{2}/\pi^{2}$ for all $x\in(0,\pi/2)$ Also it holds for all $x>\pi/2$ that
@@
$\geq\frac{2\pi\lambda^{2}k_{\textnormal{F}}^{2}}{k_{\textnormal{F}}}\sum_{k\in\Gamma}\frac{\hat{V}(k)^{2}}{2k_{\textnormal{F}}|k|}\bigg{(}\chi(2k_{\textnormal{F}}|k|t>\frac{\pi}{2})\left\{\frac{1}{2}+\log(2k_{\textnormal{F}}|k|t)-\ln(\frac{\pi}{2})\right\}$
$+\chi(2k_{\textnormal{F}}|k|t\leq\frac{\pi}{2})\frac{2(2k_{\textnormal{F}}|k|t)^{2}}{\pi^{2}}\bigg{)}\left\{1+\mathcal{O}\left(M^{\frac{1}{2}}N^{-\frac{1}{3}+\delta}+N^{-\delta}\right)\right\}$
$\eqqcolon b_{t}+d$ (6.18)
with $d\in\max\{MN^{-\frac{2}{3}+\delta},\beta k_{\textnormal{F}}^{-1}\lambda^{-1}\}$ defined in (6.10) and
$\dot{b}_{t}$ $=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{k_{\textnormal{F}}}\sum_{k\in\Gamma}\hat{V}(k)^{2}\left(\chi(2k_{\textnormal{F}}|k|t>\frac{\pi}{2})\frac{1}{2k_{\textnormal{F}}|k|t}+\chi(2k_{\textnormal{F}}|k|t\leq\frac{\pi}{2})\frac{8k_{\textnormal{F}}|k|t}{\pi^{2}}\right).$ (6.19)
@@
$g_{t}\geq b_{t}-\theta\int_{0}^{t}\textnormal{d}sg_{s}$ (6.20)
with $b_{t}$ and $\theta$ defined in (6.18) and (6.12), respectively.
@@
$h_{t}$ $=b_{t}-\theta\int_{0}^{t}\textnormal{d}sh_{s}\qquad\text{with }h_{0}=-d<0=g_{0}.$ (6.21)
@@
Assume for the proof by contradiction that there exists a $t_{0}>0:h_{t_{0}}>g_{t_{0}}$ and set $I\subseteq\mathbb{R}_{>0}$ as the largest open interval satisfying $t_{0}\in I$ and $h_{t}>g_{t}$ for all $t\in I$. Note that since $h_{t}$ is differentiable such an open interval exists. It holds $t_{1}\coloneqq\inf I>0$ since $h_{0}<g_{0}$ and for all $t\in I$
$\dot{g}_{t}\geq\dot{b}_{t}-\theta g_{t}>\dot{b}_{t}-\theta h_{t}=\dot{h}_{t}$
@@
Thus we obtain the desired contradiction to $t_{1}$ being the infimum of the largest set satisfying $h_{t}>g_{t}$. ∎
@@
$h_{t}=e^{-\theta t}\int_{0}^{t}\dot{b}_{s}e^{\theta s}\textnormal{d}s-d.$ (6.22)
@@
$h_{t}$ $=e^{-\theta t}\int_{0}^{t}\dot{b}_{s}e^{\theta s}\textnormal{d}s$
$=\frac{\pi\lambda^{2}k_{\textnormal{F}}^{2}}{k_{\textnormal{F}}}\sum_{k\in\Gamma}\hat{V}(k)^{2}e^{-\theta t}\int_{0}^{t}\bigg{(}\chi(2k_{\textnormal{F}}|k|s>\frac{\pi}{2})e^{\theta s}\frac{1}{2k_{\textnormal{F}}|k|s}$
$+\chi(2k_{\textnormal{F}}|k|s\leq\frac{\pi}{2})e^{\theta s}\frac{8k_{\textnormal{F}}|k|s}{\pi^{2}}\bigg{)}\textnormal{d}s-d$ (6.23)
$\geq\pi\lambda^{2}k_{\textnormal{F}}\sum_{k\in\Gamma}\hat{V}(k)^{2}\bigg{(}\chi(|k|t\leq\frac{\pi}{4k_{\textnormal{F}}})\frac{8k_{\textnormal{F}}|k|}{\theta^{2}\pi^{2}}\left(e^{-\theta t}+\theta t-1\right)+$
$+\chi(|k|t>\frac{\pi}{4k_{\textnormal{F}}})\frac{8k_{\textnormal{F}}|k|}{\theta^{2}\pi^{2}}\big{(}e^{-\theta\pi/(4k_{\textnormal{F}}|k|)}+\frac{\theta\pi}{4k_{\textnormal{F}}|k|}-1\big{)}\bigg{)}-d$ (6.24)
$\geq\frac{\lambda^{2}k_{\textnormal{F}}^{2}}{\theta^{2}\pi}\sum_{k\in\Gamma}\hat{V}(k)^{2}|k|\min\bigg{(}f(\theta t),f\big{(}\frac{\pi\theta}{4k_{\textnormal{F}}|k|}\big{)}\bigg{)}-d$ (6.25)
with $f(t)\coloneqq\left(e^{-t}+t-1\right)$ defining a non-negative monotonically increasing function. Recall that $d\in\max\{MN^{-\frac{2}{3}+\delta},\beta k_{\textnormal{F}}^{-1}\lambda^{-1}\}$ from (6.10) for all $0\leq t\lesssim k_{\textnormal{F}}^{-1}\lambda^{-1}$ and $\theta\leq C\lambda k_{\textnormal{F}}$ with a constant $C>0$ depending only on $V$ from (6.12). Thus, we obtain the desired result that (5.34) has a lower bound of order 1. ∎
@@ Appendix A Estimates within the bosonization framework
In this section we collect all relevant estimates of the bosonization framework which was first developed in [BNP+19, BNP+21a, BNP+21b, BPSS23] in the semiclassical regime. In addition to the references we present brief proof sketches where we think they are helpful for the interested reader.
@@
For $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta}$ and $k\in\Gamma,\alpha\in\mathcal{I}_{k}$ it holds
@@
For $N^{2\delta}\ll M\ll N^{\frac{2}{3}-2\delta}$ and $k\in\Gamma$ it holds
@@
Lemma A.3 (Estimates on the CCR error term, [BNP+21b, Lemma 5.2]).
For $k^{\prime},k\in\Gamma$ and $\alpha\in\mathcal{I}_{k},\beta\in\mathcal{I}_{k^{\prime}}$ the error term $\mathcal{E}_{\alpha}(k,k^{\prime})$ as defined in (2.22) satisfies $\mathcal{E}_{\alpha}(k,k^{\prime})=\mathcal{E}_{\alpha}(k^{\prime},k)^{\ast}$, commutes with $\mathcal{N}$ and for all $\gamma\in\mathcal{I}_{k}\cap\mathcal{I}_{k^{\prime}}$ it holds for all $\zeta\in\mathcal{F}$
@@
It holds for all $k\in\Gamma$ and $\psi\in\mathcal{F},f\in l^{2}(\mathcal{I}_{k})$:
@@
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}(k)\psi\|^{2}\leq\|\mathcal{N}^{\frac{1}{2}}\psi\|^{2}$,
@@
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}(k)\psi\|\leq M^{\frac{1}{2}}\|\mathcal{N}^{\frac{1}{2}}\psi\|$,
@@
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)\psi\|\leq M^{\frac{1}{2}}\|(\mathcal{N}+M)^{\frac{1}{2}}\psi\|$,
@@
    $\sum_{\alpha\in\mathcal{I}_{k}}\|c_{\alpha}^{\ast}(k)\psi\|^{2}\leq\|(\mathcal{N}+M)^{\frac{1}{2}}\psi\|^{2}$,
@@
    $\|\sum_{\alpha\in\mathcal{I}_{k}}f_{\alpha}c_{\alpha}^{\ast}(k)\psi\|\leq\|f\|_{l^{2}}\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|$,
@@
    $\sum_{\alpha\in\mathcal{I}_{k}}c_{\alpha}^{\ast}(k)c_{\alpha}(k)\leq\mathcal{N}$.
Lemma A.5 (Error of linearized kinetic energy, [BNP+21b, Lemma 8.2]).
It holds for $k\in\Gamma$, $\alpha\in\mathcal{I}_{k}$ and all $\psi\in\mathcal{F}$
$[\mathbb{H}_{0},c_{\alpha}^{\ast}(k)]=2k_{\textnormal{F}}|k\cdot\hat{\omega}_{\alpha}|c_{\alpha}^{\ast}(k)+\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)^{\ast}$
@@
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\psi\|^{2}$ $\leq C\left(N^{\frac{1}{3}}M^{-\frac{1}{2}}\right)^{2}\|(\mathcal{N}+1)^{\frac{1}{2}}\psi\|^{2}$
$\sum_{\alpha\in\mathcal{I}_{k}}\|\mathfrak{E}^{\textnormal{lin}}_{\alpha}(k)\psi\|$ $\leq CN^{\frac{1}{3}}\|\mathcal{N}^{\frac{1}{2}}\psi\|.$